SageG:

can someone explain to me how to find the domain of a function? especially the one with the square-roots and in a fraction as well as the ones that have x ^2. i know that it has to do with determining real numbers (i hope i am not wrong) but i dunno where to begin

SageG:

\[f(x)= \sqrt{t^2+1}\] this is an example of a question and \[g(x)=\frac{ x^2+5 }{ x+2 }\]

Hero:

For an expression such as \(\sqrt{a}\), the domain of the function is \(a \ge 0\) For an expression such as \(\dfrac{a}{b}\), the domain of the function is \(\mathbb{R}, b \ne 0\)

SageG:

thank you, can i ask a question though... in the second one, do you always just ignore the numerator no matter what the number is?

Hero:

False. If it has a square root in the numerator, for example, you still have to consider it.

Hero:

\(f(x) = \sqrt{\dfrac{2x}{x-2}}\) What is the domain of this for example?

SageG:

okay i get it. another question please, does the addition and subtraction matter in the calculations? someone once told me that because there is an addition in the square root of a function, it can be ignored

Hero:

Example of what you're referring to please.

Hero:

NO such thing as ignoring without proper justification in mathematics

SageG:

okay... so it was something like \[\frac{ x+4 }{ \sqrt{x-4} }\] and so the person solved the one above and said that you cannot solve the one denominator because there is a minus in it. im sorry for the late reply, i was looking for the question but i couldnt find it so i improvised instead

Hero:

You should probably find the question.

Hero:

If the question is to find the domain, then the domain will be \(\sqrt{x-4} > 0\)

Hero:

From there you solve for \(x\)

Hero:

@SageG any questions concerning that?

SageG:

alright i will seach for it. and yes i have a question, regarding solving x, is the first step separating the square root into two parts, i.e. square root x minus square root of 4 ?

Hero:

There's no such rule as what you are describing. You're asking if \(\sqrt{x - a} = \sqrt{x} - \sqrt{a}\) . Where have you ever seen a rule such as that?

SageG:

i havent, my sister did that and said she isnt sure so i should ask

Hero:

To solve \(\sqrt{x - 4} > 0\) , you begin by squaring both sides.

Hero:

\((\sqrt{x - 4})^2 > 0^2\)

SageG:

why?

Hero:

Because the goal is to solve for \(x\)

Hero:

In order to solve for \(x\) you can only use the rules of mathematics to do so.

Hero:

In order to isolate \(x\) in this case, you must get rid of the square root. The only way to do that is to perform the inverse operation and apply it to both sides of the equation. The inverse of square root is square.

SageG:

okay so, i understand. then x-4>0 which in turn will become x>4

Hero:

Correct

SageG:

oh okay, thank goodness. okay so back to your question f(x) = square root 2x / x-2 x-2 \[x-2\neq0 \] that is to say as long as x is not 2, the function has a defined value. then\[2x \ge0\] ....is that wrong?

TETSXPREME:

*welcome to QC. we're excited to have you with us!*